OpenStudy (anonymous):

Simplify. 4 to the power of 2 + 3 to the power of 3 A. 17 B. 35 C. 43 D. 97 @imqwerty

OpenStudy (anonymous):

43.

OpenStudy (anonymous):

4^2 = 16 3^3 = 27 16+27= 43

OpenStudy (anonymous):

Thx @xFrenchFryx

OpenStudy (anonymous):

Np. :3

OpenStudy (danjs):

4*4 + 3*3*3